Using the discriminant, determine the number of x-intercepts the graph of the equation will have. Do not graph the equation.
y = -5x^2 + 4x - 1

Answer by Alan3354(69443)

y = -5x^2 + 4x - 1
--------
a = -5
b = +4
c = -1
------
Discriminant = b^2 - 4ac = 16 - 20
Disc <0 --> no real solutions, no crossing of the x-axis.
-------------
I might have graphed it, but you can't prove it.